OpenStudy (anonymous):

whats the answer too 6+9+66-2332+00980000-11112+3

Nnesha (nnesha):

fist add 6 9 3 and 66

Nnesha (nnesha):

9 + 6 = 15 15 + 3 = 18 18 + 66 = 84 84 +00980000= 980084 980084 -11112=968972 968972 - 2332= 966640

OpenStudy (anonymous):

thanks

Nnesha (nnesha):

\(\huge\color{black}{my ~ pleasure }\)